One angle of a triangle measures 50 degrees. The other two angles are in a ratio of 6:7. What are the measures of those two angles?
Answer by ikleyn(52824)

.

The two unknown angles are 6x and 7x  degrees.


The sum of the angles equation is


    50 + 6x + 7x = 180  degrees


           13x   = 180 - 50

           13x   = 130 degrees

             x   = 130/13 = 10 degrees.


ANSWER.  The two unknown angles are 60 degrees and 70 degrees.

Solved.